OpenStudy (anonymous):

Suppose the population of a town is 8,600 and is growing 3% each year. Predict the population after 3 years.

OpenStudy (anonymous):

\[8600 \times \left(1+\frac 3 {100}\right)^3 \]

Parth (parthkohli):

= 8600 + 258 + 265 + 273 = 9 396

OpenStudy (anonymous):

Suppose that the population of deer in a state is 1,500 and is growing 2% each year. Predict the population after 4 years.